By what amount does the 52-cm-long femur of an 73 kg runner compress at this moment? The cross-section area of the bone of the femur can be taken as 5.2×10?4m2 and its Young's modulus is 1.6×1010N/m2.
Answer #1
Answer #2

Young's Modulus is given as

Y=\frac{\frac{F}{A}}{\frac{\Delta L}{L}}

Therefore, change in length is given as

\Delta L=\frac{F\times L}{A\times Y}\Rightarrow \frac{73 \times 9.8 \times 0.52 }{5.2\times 10^{-4}\times 1.6\times 10^{10}}=44.71 \times 10^{-6} \ m \ i.e \ 44.71 \ \mu m